# Orbis: Overcoming Challenges of Long-Horizon Prediction in Driving World Models

### v1

@@ Abstract
Existing world models for autonomous driving struggle with long-horizon generation and generalization to challenging scenarios. In this work, we develop a model using simple design choices, and without additional supervision or sensors, such as maps, depth, or multiple cameras. We show that our model yields state-of-the-art performance, despite having only 469M parameters and being trained on 280h of video data. It particularly stands out in difficult scenarios like turning maneuvers and urban traffic. We test whether discrete token models possibly have advantages over continuous models based on flow matching. To this end, we set up a hybrid tokenizer that is compatible with both approaches and allows for a side-by-side comparison. Our study concludes in favor of the continuous autoregressive model, which is less brittle on individual design choices and more powerful than the model built on discrete tokens. Code, models and qualitative results are publicly available at https://lmb-freiburg.github.io/orbis.github.io/.
@@ 1 Introduction
Intelligent agents operate in complex environments by simulating plausible future states based on past observations. This capacity for imagination allows them to plan toward long-term goals bar2024navigation ; ha2018world . Humans naturally acquire this ability through passive observation and minimal interactions, enabling them to adapt quickly to new and unseen scenarios. Emphasizing the passive observation component of such world models has become particularly popular for the driving world, since large amounts of data exists for this domain, and it is attractive to circumvent the manual setup of many perception components by rather learning the visual representation for decision making via the predictive loss of a world model.
@@
Recent driving world models agarwal2025cosmos ; gao2024vista ; hassan2024gem built on video diffusion models blattmann2023stable have made major strides towards generating detailed content in high definition and at high frame-rates. However, Fig. 2 reveals that these models only work well for few frames, especially in case of maneuvers that require generating new content, such as turning. This indicates limitations in how these models capture state transitions – the key feature of a world model. We quantified this shortcoming via a metric that compares the trajectory of the real continuation with the trajectories of the generated content. Our evaluation reveals consistent problems with all public driving world models.
@@
Consequently, a relevant question is whether world models should rely on continuous-space modeling or predict discrete tokens (similar to LLMs)chang2022maskgit ; esser2021taming ; yu2023magvit . The current trend for visual generation goes towards diffusion-based (continuous) models blattmann2023stable ; polyak2024movie ; girdhar2024factorizing . On the other hand, driving world models based on discrete representations and LLM-like objectives seem to have the edge in terms of rollout duration hu2023gaia ; hu2024drivingworld . Among these, the proprietary GAIA-1 model showed no issues with turns and long drives. This observation prompts the question of whether the discrete paradigm is really superior to the other for long-term generation, and whether the continuous space is the reason for the observed shortcomings in the current state of the art.
@@
Figure 1: Comparison of model scale, training data volume, and FVD performance of various approaches on the NuPlan dataset. *DrivingWorld is trained on the test dataset nuPlan.
@@
To summarize, (1) we highlight shortcomings of contemporary driving world models and propose additional benchmarking metrics to make these shortcomings more explicit. (2) We propose a hybrid discrete-continuous tokenizer that is compatible with both discrete and continuous prediction losses and allows us to compare both modeling approaches side-by-side. (3) On its basis, we compare continuous and discrete prediction losses on a fair common ground and find a clear advantage in favor of continuous modeling. (4) As demonstrated in Figure 1, the resulting model is much more economical in terms of training data and model size than existing world models. Using only 280 hours of front camera video data, our 469M parameter flow-matching model Orbis already produces state-of-the-art performance on long-horizon rollouts with realistic and diverse trajectories. It excels particularly in challenging driving scenarios. Find visualizations embedded in the attached html file.
@@ 2 Related Work
World Models. The ability of world models to do real-world simulation can be useful for policy learning he2024large ; piergiovanni2019learning , sample efficient RL hafner2019dream ; hafner2020mastering ; li2024think2drive ; wu2023pre , and representation learning zhao2024drivedreamer4d . Previous world models have been limited to gaming ha2018recurrent ; hafner2019dream ; hafner2020mastering and other simulated environments dosovitskiy2017carla . Recent breakthroughs in video generative modeling blattmann2023stable ; yang2024cogvideox have led to future video prediction models - an essential building block for world models.
@@
Figure 2: Limitations of state-of-the-art video generation models on turning events. Left: The trajectories estimated from the generated videos show that previous approaches either stop prematurely or drift into an unnatural path. Right: The quality of the corresponding generated frames degrades over time, as the models struggle to generate the scenery. In contrast, our method tracks the original trajectory curvature and speed more closely, and can generate novel content beyond the unseen horizon.
@@ 3 Challenges Faced by Contemporary Driving World Models
These are not well captured by standard metrics like FVD, whose reliance on pretrained human-action encoders limits sensitivity to scene dynamics and ego-motion. This calls for more targeted evaluations. To bridge this gap, we propose a distribution-level trajectory-based evaluation, detailed in Section 5.2.2, that directly quantifies realism and coverage of generated driving behavior compared to a curated dataset of turning events. We evaluate and compare the generated trajectories for Vista, GEM, and our approach, and find the results to confirm our qualitative observations and show the shortcomings of the existing methods.
@@ 4.2.1 Flow Matching
$\mathcal{L}=\mathbb{E}_{\tau\sim[0,1],\,\epsilon\sim\mathcal{N}(0,I)}\big{[}||\mathbf{v}(\mathbf{x}_{N}^{\tau};\mathbf{x}_{0:N-1})-(\epsilon-\mathbf{x}_{N})||^{2}\big{]},$ (2)
@@ 5.1 Experiment Details
To train our world model, we use subsets of videos from the BDD100K yu2020bdd100k and OpenDV yang2024genad datasets. As shown in Table 1, we select a limited number of hours from each dataset and extract frames at 10 Hz. In total, we use 280 hours of video data from a combined available total of 2747 hours. For BDD100K, we select the day-clear subset of the training set. From OpenDV we exclude night drives via a brightness filter and uneventful ones by the presence of certain words in the original video titles (see Appendix B). We then subsample by selecting evenly spaced 30-second clips. To train the tokenizer we additionally select images from Honda HAD kim2019CVPR , Honda HDD ramanishka2018CVPR , ONCE mao2021one , NuScenes Caesar_2020_CVPR , and NuPlan caesar2021nuplan to make the dataset diverse. Our dataset primarily consists of daylight scenarios.
@@
For the FM model, we use DiT peebles2023scalable with ST transformer blocks (STDiT). To improve generalization and frame generation quality, we drop all context frames 50% of the time. When context frames are present, we augment them with noise 50% of the time, similar to prior work valevski2024diffusion ; gao2025adaworld ; he2022latent . In order to sample the next frame, we use ODE sampler and take 30 steps. For the MGM model, we also add context noise to improve robustness towards context noise, especially for long-rollouts: we replace 10% of the frames and 10% of the overall tokens with a mask token.
@@
Training details. Ours high-resolution model operates at $512\times 288$ and small-scale model at $256\times 256$. Tokenizer compresses the image spatially by 16×16$\times$16 ×. We train latent models with a context of 5 frames sampled at 5Hz. All small-scale models for ablation studies are trained on only the BDD100K subset for one day on 32×32$\times$32 × A100 GPUs. The high resolution model is trained for 10 epochs over 5 days on 72×72$\times$72 × A100 GPUs.
Table 2: SOTA results: FVD over 6 seconds rollouts@ 5Hz. Numbers of baselines were computed using their official checkpoints. Lower FVD is better. *DrivingWorld (DW) is trained on the test dataset nuPlan and uses ego-motion control as an extra input. Video samples available on the project page.
Model nuPlan Waymo nuPlan
turns
Cosmos agarwal2025cosmos 291.80 278.19 316.99
Vista gao2024vista 323.37 422.58 413.61
GEM hassan2024gem 431.69 291.84 357.25
DW* hu2024drivingworld 298.97 N/A 334.89
Orbis (ours) 132.25 180.54 231.88
@@ 5.2.1 Video Generation Quality
We evaluate the quality of the generated videos using FVD  unterthiner2019fvd and FID. For comparing with the baselines, we show FVD on nuPlan caesar2021nuplan and Waymo sun2020scalability datasets with 800 and 400 samples respectively. To evaluate the models in challenging scenarios, we curate a dedicated validation set of turning events (nuPlan-turns), consisting of 400 samples, selected from the nuPlan validation set where the initial yaw rate is at least 12rad/s ($\sim$1 std). FVD results on nuPlan and nuPlan-turns are not comparable, since nuPlan is a much more diverse dataset compared to the specialized nuPlan-turns dataset. All evaluation datasets constitute unseen testing domains for our model and for the baselines, except for DrivingWorld, which contains nuPlan as part of the training dataset. We choose nuPlan over the similar nuScenes due to the latter’s irregular sampling rate, which adds an unnecessary confounding factor to the evaluation.
@@ 5.2.2 Trajectory Quality
We compare the quality of generated trajectories for Vista, GEM, and our approach in (Table 3). The results show the limitations of existing models in capturing the underlying distribution of ego motion and driving behavior. All approaches perform worse in terms of ADE, indicating difficulty in maintaining realistic speeds. Moreover, we achieve the best precision-recall on Fréchet distance, indicating that our predicted trajectories more closely follow the ground-truth paths compared to competing baselines.
@@
$+$ ego-motion 2.40
5.3 Results
Comparison to SOTA.
Results are shown in the Table 2, for a prediction horizon of 6 seconds at 5hz. The qualitative results for 20s are shown in Figure 5( more are included on the project page). Our method outperforms other driving world models on all three benchmarks. We further compare results for long-horizon video prediction, shown in Figure 4 on nuPlan-turns dataset. For each method, FVD is computed over the entire predicted video in a windowed manner, where each window contains 20 frames sampled at 5hz. Results show that Orbis based on flow matching outperforms all baselines and maintains stable performance over long-horizon prediction of up to 20 seconds. The discrete counterpart Orbis-MG based on MaskGIT, shows suboptimal performance for shorter horizons but scales well over long horizons, surpassing all previous works for the last two windows. As discussed earlier, previous works perform well in short horizons but struggle with long-horizon predictions. GEM has higher FVD scores for short-horizon due to its single-frame conditioning design but performs relatively better on long-horizon predictions (more details in the Appendix E).
Ego-motion Control and Evaluation
As a proof-of-concept for ego-motion control, we fine-tune a copy of high resolution Orbis for two epochs on 75h of nuPlan videos and IMU data. Following previous literature gao2024vista ; hassan2024gem ; hu2024drivingworld , we evaluate the resulting model by computing the ADE pellegrini2009you between true and generated trajectories, estimated with VGGT. We compare the ADE of the same model with and without steering on 400 5s long nuPlan validation sequences in Table 5. Better trajectory tracking under ego-motion conditioning indicates some degree of controllability – though in a preliminary setting. Indeed, conditioning capabilities for related models are well documented peebles2023scalable ; bar2024navigation .
Effect of tokenizer design
For the discrete model, adding DINO distillation to the image tokenizer, similar to GAIA-I hu2023gaia , leads to lower FVD, as shown in Table 5. However, the key factor to enable long-horizon prediction for the discrete model is token factorization. Usage of DINO distillation even leads to a worse rFID (reconstruction FID). However, token factorization annihilates this difference. Interestingly, while the factorized tokenizer with DINO distillation is very important for the discrete model, the continuous model is robust to these design changes, showing no large change in FVD. These experiments were conducted in the small-scale setting.
@@ 5.3 Results
Figure 5: Qualitative results of the Orbis model over 20-sec rollouts (zoom-in for details). Videos samples are available on the project page.

### v2

@@ Abstract
Existing world models for autonomous driving struggle with long-horizon generation and generalization to challenging scenarios. In this work, we develop a model using simple design choices, and without additional supervision or sensors, such as maps, depth, or multiple cameras. We show that our model yields state-of-the-art performance, despite having only 469M parameters and being trained on 280h of video data. It particularly stands out in difficult scenarios like turning maneuvers and urban traffic. We test whether discrete token models possibly have advantages over continuous models based on flow matching. To this end, we set up a hybrid tokenizer that is compatible with both approaches and allows for a side-by-side comparison. Our study concludes in favor of the continuous autoregressive model, which is less brittle on individual design choices and more powerful than the model built on discrete tokens. Project page with code, model checkpoints and visualization can be found here: https://lmb-freiburg.github.io/orbis.github.io
@@
{mousakha,mittal,galessos,faridk,brox}@cs.uni-freiburg.de
@@ 1 Introduction
Intelligent agents operate in complex environments by simulating plausible future states based on past observations. This capacity for imagination allows them to plan toward long-term goals ha2018world; bar2024navigation. Humans naturally acquire this ability through passive observation and minimal interactions, enabling them to adapt quickly to new and unseen scenarios. Emphasizing the passive observation component of such world models has become particularly popular for the driving world, since large amounts of data exists for this domain. It is attractive to circumvent the manual setup of many perception components by learning the visual representation for decision making via the predictive loss of a world model.
@@
Recent driving world models gao2024vista; agarwal2025cosmos; hassan2024gem built on video diffusion models blattmann2023stable have made major strides towards generating detailed content in high definition and at high frame-rates. However, Fig. 2 highlights that these models only work well for few frames, especially in case of maneuvers that require generating new content, such as turning. Our evaluation, conducted on the generated videos as well as on the estimated ego-trajectories, reveals substantial limitations in how all public driving world models capture state transitions – the key feature of a world model – even though the quality of the generated videos is excellent at first.
While realistic video generation can be a valuable product, it is not the primary objective of world models. As demonstrated in a number of influential works (e.g. World Models ha2018world, Dreamer hafner2019dream, Cosmos agarwal2025cosmos, VJEPA-2 assran2025v), their main purpose is representation learning, planning, and policy learning. The quality of state representations and the accuracy of next-state predictions are therefore paramount, with decoded videos serving primarily as an indicator, e.g. of whether the model can execute a turn stably. Long-horizon prediction serves as a more meaningful measure of how well the model captures state transitions, while generalization to complex scenarios reflects its ability to model diverse real-world dynamics. Existing models such as Vista and GEM, though conceived with planning as possible application, often fail in non-trivial yet routine situations. In this work we prioritize building a world model with robust state representation and dynamics that handle such cases effectively.
@@
Consequently, a relevant question is whether world models should rely on continuous-space latents or use discrete tokens (similar to LLMs)esser2021taming; chang2022maskgit; yu2023magvit for representing world states. The current trend for visual generation goes towards diffusion-based (continuous) models polyak2024movie; blattmann2023stable; girdhar2024factorizing. On the other hand, driving world models based on discrete representations and LLM-like objectives seem to have the edge in terms of rollout duration hu2023gaia; hu2024drivingworld. Among these, the proprietary GAIA-1 model showed no issues with turns and long drives. This observation prompts the question of whether the discrete paradigm is really superior to the other for long-term generation, and whether the continuous space is the reason for the observed shortcomings in the current state of the art.
@@
Figure 1: Comparison of model scale, training data volume, and FVD performance of various approaches on the NuPlan-turns dataset. *DrivingWorld is trained on the test dataset nuPlan.
@@
To summarize, (1) we highlight shortcomings of contemporary driving world models and propose additional benchmarking metrics to make these shortcomings more explicit. (2) We propose a hybrid discrete-continuous tokenizer that is compatible with both discrete and continuous prediction losses and allows us to compare both modeling approaches side-by-side. (3) On its basis, we compare continuous and discrete prediction losses on a fair common ground and find a clear advantage in favor of continuous modeling. (4) As demonstrated in Figure 1, the resulting model is much more economical in terms of training data and model size than existing world models. Using only 280 hours of front camera video data, our 469M parameter flow-matching model Orbis already produces state-of-the-art performance on long-horizon rollouts with realistic and diverse trajectories. It excels particularly in challenging driving scenarios.
@@ 2 Related Work
World Models. The ability of world models to do real-world simulation can be useful for policy learning piergiovanni2019learning; he2024large, sample efficient RL hafner2019dream; hafner2020mastering; wu2023pre; li2024think2drive, and representation learning zhao2024drivedreamer4d. Previous world models have been limited to gaming hafner2020mastering; hafner2019dream; ha2018recurrent and other simulated environments dosovitskiy2017carla. Recent breakthroughs in video generative modeling yang2024cogvideox; blattmann2023stable have led to future video prediction models - an essential building block for world models.
@@
Figure 2: Limitations of state-of-the-art video generation models on turning events. Left: The trajectories estimated from the generated videos show that previous approaches either stop prematurely or drift into an unnatural path. Right: The quality of the corresponding generated frames degrades over time, as the models struggle to generate the scenery. In contrast, our method tracks the original trajectory curvature and speed more closely, and can generate novel content beyond the unseen horizon. Videos for a larger set of randomly sampled context frames are linked in the Appendix.
@@ 3 Challenges Faced by Contemporary Driving World Models
These are not well captured by standard metrics like FVD or JEDi, whose reliance on pretrained human-action or general purpose encoders limits sensitivity to driving dynamics and ego-motion. This calls for more targeted evaluations. To bridge this gap, we propose a distribution-level trajectory-based evaluation, detailed in Section 5.2.2, that directly quantifies realism and coverage of generated driving behavior compared to a curated dataset of turning events. We evaluate and compare the generated trajectories for Vista, GEM, and our approach, and find the results to confirm our qualitative observations and show the shortcomings of the existing methods.
@@ 4.2.1 Flow Matching
[…]
@@ 5.1 Experiment Details
To train our world model, we use subsets of videos from the BDD100K yu2020bdd100k and OpenDV yang2024genad datasets. As shown in Table 1, we select a limited number of hours from each dataset and extract frames at 10 Hz. In total, we use 280 hours of video data from a combined available total of 2747 hours. For BDD100K, we select the day-clear subset of the training set. From OpenDV we exclude night drives via a brightness filter and uneventful ones by the presence of certain words in the original video titles (see Appendix). We then subsample by selecting evenly spaced 30-second clips. To train the tokenizer we additionally select images from Honda HAD kim2019CVPR, Honda HDD ramanishka2018CVPR, ONCE mao2021one, NuScenes Caesar_2020_CVPR, and NuPlan caesar2021nuplan to make the dataset diverse. Our dataset primarily consists of daylight scenarios.
@@
For the FM model, we use DiT peebles2023scalable with ST transformer blocks (STDiT), where a per-frame causal attention mask is applied to the time-attention layers. To improve generalization and frame generation quality, we drop all context frames 50% of the time. When context frames are present, we augment them with noise 50% of the time, similar to prior work valevski2024diffusion; gao2025adaworld; he2022latent. In order to sample the next frame, we use ODE sampler and take 30 steps. For the MGM model, we also add context noise to improve robustness towards context noise, especially for long-rollouts: we replace 10% of the frames and 10% of the overall tokens with a mask token.
@@
Training details. Our higher resolution model operates at $512\times 288$ and small-scale model at $256\times 256$. Tokenizer compresses the image spatially by $16\times$. We train latent models with a context of 5 frames sampled at 5Hz. All small-scale models for ablation studies are trained on only the BDD100K subset for one day on $32\times$ A100 GPUs. The higher resolution model is trained for 10 epochs over 5 days on $72\times$ A100 GPUs.
@@ 5.2.1 Video Generation Quality
We evaluate the quality of the generated videos using FVD  unterthiner2019fvd and JEDi luo2024jedi, and we use FID to assess tokenizer reconstruction quality. For comparing with the baselines, we show FVD and JEDi on nuPlan caesar2021nuplan and Waymo sun2020scalability datasets with 800 and 400 samples respectively. To evaluate the models in challenging scenarios, we curate a dedicated validation set of turning events (nuPlan-turns), consisting of 400 samples, selected from the nuPlan validation set where the initial yaw rate is at least 0.12 rad/s ($\sim$1 std). FVD results on nuPlan and nuPlan-turns are not comparable, since nuPlan is a much more diverse dataset compared to the specialized nuPlan-turns dataset. All evaluation datasets constitute unseen testing domains for our model and for the baselines, except for DrivingWorld, which contains nuPlan as part of the training dataset. We choose nuPlan over the similar nuScenes due to the latter’s irregular sampling rate, which adds an unnecessary confounding factor to the evaluation. Additionally, we evaluate our model using Video Quality Assessment (VQA) metrics including PSNR, SSIM and DOVERwu2023exploring, see Appendix D.2.
@@ 5.2.2 Trajectory Quality
We compare the quality of generated trajectories for Vista, GEM, and our approach in (Table 2). The results show the limitations of existing models in capturing the underlying distribution of ego motion and driving behavior. All approaches perform worse in terms of ADE, indicating difficulty in maintaining realistic speeds. Moreover, we achieve the best precision-recall on Fréchet distance, indicating that our predicted trajectories more closely follow the ground‐truth paths compared to competing baselines.
5.3 Results
Table 3: SOTA results: FVD and JEDi over 6 seconds rollouts@ 5Hz. Numbers of baselines were computed using their official checkpoints. Lower is better. *DrivingWorld (DW) is trained on the test dataset nuPlan and uses ego-motion control as an extra input. Video samples available in Appendix.
[…]
Model nuPlan Waymo nuPlan nuPlan Waymo nuPlan
turns turns
Cosmos agarwal2025cosmos 291.80 278.19 248.39 0.55 0.31 0.50
Vista gao2024vista 323.37 422.58 413.61 0.37 0.44 0.54
GEM hassan2024gem 431.69 291.84 357.25 0.42 0.35 0.31
DW* hu2024drivingworld 298.97 N/A 334.89 1.33 N/A 1.50
Orbis (ours) 134.06 167.57 239.20 0.14 0.33 0.16
Comparison to SOTA.
Results are shown in the Table 3, for a prediction horizon of 6 seconds at 5hz. The qualitative results for 20s are shown in Figure 5 (more are included in Appendix). Our method outperforms other driving world models on all three benchmarks and both metrics, except for Cosmos on Waymo according to the JEDi metric. We further compare results for long-horizon video prediction, shown in Figure 4 on nuPlan-turns dataset. For each method, FVD is computed over the entire predicted video in a windowed manner, where each window contains 20 frames sampled at 5hz. Results show that Orbis based on flow matching outperforms all baselines and maintains stable performance over long-horizon prediction of up to 20 seconds. The discrete counterpart Orbis-MG based on MaskGIT, shows suboptimal performance for shorter horizons but scales well over long horizons, surpassing all previous works for the last two windows. As discussed earlier, previous works perform well in short horizons but struggle with long-horizon predictions. GEM has higher FVD scores for short-horizon due to its single-frame conditioning design but performs relatively better on long-horizon predictions (more details in the Appendix E).
Ego-motion Control and Evaluation
As a proof-of-concept for ego-motion control, we fine-tune a copy of Orbis model for two epochs on 75h of nuPlan videos and IMU data. Following previous literature gao2024vista; hassan2024gem; hu2024drivingworld, we evaluate the resulting model by computing the ADE pellegrini2009you between true and generated trajectories, estimated with VGGT. We compare the ADE of the same model with and without steering on 400 5s long nuPlan validation sequences in Table 5. Better trajectory tracking under ego-motion conditioning indicates some degree of controllability – though in a preliminary setting. Indeed, conditioning capabilities for related models are well documented peebles2023scalable; bar2024navigation.
@@ 5.3 Results
++ ego-motion 2.40
Effect of tokenizer design.
For the discrete model, adding DINO distillation to the image tokenizer, similar to GAIA-1 hu2023gaia, leads to lower FVD, as shown in Table 5. However, the key factor to enable long-horizon prediction for the discrete model is token factorization. Usage of DINO distillation even leads to a worse rFID (reconstruction FID). However, token factorization annihilates this difference. Interestingly, while the factorized tokenizer with DINO distillation is very important for the discrete model, the continuous model is robust to these design changes, showing no large change in FVD. These experiments were conducted in the small-scale setting.
@@
Table 6: FVD scores for Orbis-MG and Orbis-FM model with different architectures. The FVD is over 60 generated frames of BDD val set.
Model Architecture 12 steps 30 steps
Orbis-MG ST 533.3 571.7
(discrete) DiT 769.0 981.3
CDiT 1552.3 1718.9
Orbis-FM ST 360.9 246.0
(conti.) DiT 345.4 274.2
CDiT 410.2 246.1
Consistency over architectures and sampling budgets.
To assess the generality of our findings, we evaluate Orbis-FM and MG using three Transformer architectures – DiT peebles2023scalable, STDiT (our default), and CDiT bar2025navigation, and with different sampling budgets. All models were trained for one day on 32 A100 GPUs in a small-scale setting, and FVD scores were computed on 200 generated BDD validation videos (60 frames at 5 fps). As shown in Table 6, Orbis-FM consistently outperforms Orbis-MG across all architectures and inference settings (12 and 30 steps). Orbis-FM shows greater consistency across architectures than Orbis-MG. The ST architecture shows the best results for both models, with Orbis-FM (ST, 30 steps) achieving the best performance overall.
Figure 5: Qualitative results of the Orbis model over 20-sec rollouts (zoom-in for details). Videos and more samples available on the webpage.